# Argumentatively Coherent Judgmental Forecasting

### v1

@@ 6.2.2 ArguCast with ArgLLMs
In order to populate ArguCast, we used LLMs to get predictions and ArgLLMs, leveraging LLMs, to construct QBAFs (we used the same prompts and hyperparameters as in Freedman et al. (2025)): we prompted the LLMs to generate supporting and attacking arguments; for each generated argument, we prompted the LLMs again to obtain an uncertainty score (acting as the base score). We investigated two variations of ArgLLMs: generating arguments at breadth $,1$ 1 (only generating a supporter and an attacker for each forecasting argument), and breadth $n,k$ (generating $n$ supporting and $k$ attacking arguments for each forecasting argument). We set the base score for each forecasting argument to 0.5. Figure 3 shows an example (at breadth $,1$ 1).
@@
Figure 3: An example of an ArgLLM (with Mistral-7B-Instruct-v0.3 as the underpinning LLM), applied to the forecasting argument $f=$‘SpaceX’s next Starship orbital flight test will not be launched successfully before 15 October 2021’, generating a QBAF with a supporting and an attacking argument, with base scores 0.6 and 0.4 respectively, leading to a strength $\sigma(f)=0.6$ with Df-QuAD (given base score $\tau(f)=0.5$ ). The resulting forecaster QBAF leads to an incoherent forecast, as the initial prediction, 40%, is lower than $\xi_{2}(f)=0.5$, but $\sigma(f)>\xi_{1}(f)=0.5$.
@@ 6.3 Results
Table 2 reports the forecasting accuracy for each LLM and the forecasting accuracy in the set of coherent predictions for both breadth $,1$ 1 and breadth $n,k$. The results indicate that applying coherence generally leads to improvements or, at least, preserves forecasting accuracy across all models. Note that applying coherence significantly reduces the total number of predictions, at breadth $n,k$ generally retaining more samples compared to breadth $n,k$. Most notably, Llama 3 retains only 44% of samples at breadth $,1$ 1, while it preserves 71% at breadth $n,k$. The most samples retained were with the Mistral model at breadth $n,k$, preserving 78% of the samples.
@@
Table 2: Accuracy comparison between raw LLMs and ArgLLMs with breadth $,1$ 1 and breadth $n,k$. For raw LLMs, we show accuracy as (correct/total). For ArgLLMs we show the percentages of coherent and correct predictions within those of coherent predictions (coherent & correct/total coherent). The bold percentage shows the best variant for the given model. The total columns show the number of samples retained for each model.
The key finding is that accuracy consistently improves when coherence is applied, particularly at breadth $,1$ 1. GPT-4o shows the most dramatic improvement, with accuracy increasing from 69% to 80% at breadth $,1$ 1. Llama 3 and Mixtral also achieve their best performance when coherence is applied at breadth $,1$ 1, reaching 82% and 78% accuracy, respectively. Interestingly, Mistral is the only model that achieved a higher accuracy at breadth $n,k$ compared to breadth $,1$ 1, which could suggest that already better performing models may require more arguments to be generated.
These results support a positive answer to RQ2. Notably, the improvements are most consistent at breadth $,1$ 1.
@@ 7 Crowd-sourced Experiments
Figure 4: Example of a debate shown to users. The debate includes a forecasting argument (outlined in cyan), a supporter thereof (outlined in gray with a green solid arrow pointing towards the forecasting argument), an attacker of the forecasting argument (also outlined in gray but with a red dotted arrow pointing towards the forecasting argument), and the votes of a fictitious forecaster $u$ (where blue thumbs up represents $V(u,a)=+$ and pink thumbs down represents $V(u,a)=-$).
@@ 7.1 Research Questions
We anticipated that, for RQ3, the null hypothesis (stating that users’ reasoning and our notion of coherence disagree in the same manner) would not be rejected. However, for RQ4, we expected the null hypothesis to be rejected (i.e., that RQ4 would be answered affirmatively).
@@
$|\mathcal{F}|=1\land|\mathcal{D}|=2$ $\land|\mathcal{A}|=1\land|\mathcal{S}|=1\land$
$(a,f)\in\mathcal{A}\rightarrow\mathcal{V}(u,a)=+$ $\land(s,f)\in\mathcal{S}\rightarrow\mathcal{V}(u,s)=-$
@@
$(\exists_{c\in\mathcal{D}}\exists_{d\in\mathcal{D}}\exists_{g\in\mathcal{F}}(c,g)\in\mathcal{A}$ $\land(d,g)\in\mathcal{S}$
$\land\mathcal{V}(u,c)=\mathcal{V}(u,d))$
@@
A breadth complex variant is an ACF $\langle\mathcal{X}=\mathcal{F}\cup\mathcal{D},\mathcal{R}=\mathcal{A}\cup\mathcal{S},\mathcal{U},\mathcal{V},\mathcal{P}\rangle$, such that
|$\lvert\{a\in\mathcal{D}|\not\exists_{r\in\mathcal{R}}{(r=(a_{1},a),a_{1}\in\mathcal{D})}\}\rvert=3$
@@
|$|\{a\in\mathcal{D}|\exists_{r\in\mathcal{R}}{(r=(a,a_{1}),a_{1}\in\mathcal{X})}\}|=1$
A depth complex variant occurs when the forecaster adds a new argument to any of the leaf (i.e. unattacked and unsupported) nodes. This argument can either support or attack one of the child arguments (the supporter or attacker of the forecasting argument).9
@@ 7.2 Experimental Design
[…]
@@ 7.4 Results
[…]
@@
Table 3: Contingency table for McNemar’s test. Columns represent the number of debate variants the users found coherent and rows represent the number of debate variants our notion of coherence found coherent for the same debate variants.
@@
$\neg$¬Aligned 26 9 6 8 9 3 8 19
Table 4: Number of aligned and not-aligned pairs (alignment between our notion of coherence and the participants in the experiment) for the debate variations: simple, breadth complex, depth complex and vote complex. The mean of the aligned pairs used in the statistical tests are calculated from these numbers.
Table 5 summarises the results for RQ4.1-RQ4.3. For each complexity type, we conducted a two-sample one sided t-test to determine if complex debate variants aligned more with our notion of coherence compared to the simpler variants. Our null hypothesis for all RQ4.1-RQ4.3 was H0: $\mu_{\text{complex}}\leq\mu_{\neg\text{complex}}$ against the alternative H1: $\mu_{\text{complex}}>\mu_{\neg\text{complex}}$. All three complexity types were statistically significant at $p<0.05$. Vote complexity yielded a p-value of 0.009, breadth complexity a p-value of 0.04, and depth complexity the most significant effect with a p-value of 0.002. Based on these results, we can reject all three null hypotheses in favor of their alternatives.
@@
Table 5: This table shows the mean and standard deviation of aligned pairs (between our notion of coherence and the participants from the user experiment) for each complex and not-complex debate variations. These values are derived from Table 4. Then, it reports on the p-value derived from the two-sample one sided t-test.
Rejecting RQ3 suggests that the notion of coherence with $\xi_{1}(f)=\xi_{2}(f)=0.5$ does not accurately reflect the intuition of what users perceive as coherent. This further implies the need for a coherence constraint, as users are not naturally coherent in their judgment. Additionally, since we have rejected RQ4.1, RQ4.2, and RQ4.3 and these were refinements of RQ4, we can also reject RQ4. This means that our notion of coherence aligns more with users’ reasoning when debates are complex. This could suggest that our notion of coherence is more intuitive when more information is available to the users. Furthermore, this indicates that with complex debate variations, the users spend more cognitive effort with complex debate variations, which helps them counteract their biases.
@@ 8 Conclusion
In this paper, we introduced a notion of argumentative coherence for judgmental forecasting, formally defining it in terms of alignment between forecasters’ internal argumentative reasoning and their predictions. Then, through crowd-sourced and in-person user studies, as well as experiments with LLMs, we systematically evaluated the usefulness of the coherence property.
First, in an in-person user study with the ArguCast system, we found that filtering out incoherent forecasts improved the aggregated predictions, moving them closer to expert baselines. Second, in the experiments with LLMs, we applied coherence to LLM-generated forecasts and ArgLLM-generated argumentative reasoning and demonstrated consistent improvements in accuracy across different models. These results from the in-person and LLM experiments support the practical need for integrating argumentative coherence into judgmental forecasting systems (like ArguCast, GJOpen, and Metaculus). Finally, in the crowd-sourced experiments, we found that the users’ opinions may not align with (our notion of) coherence. This suggests that some form of coherence constraints are necessary. Additionally, we observe that the users’ opinions aligned more with our notion of coherence when debate variants were more complex. We speculate that this is probably due to our notion being more intuitive when more information is available to the participants. The results also indicated that as users spend more cognitive effort with complex debate variations (which implies that they reason more), it helps them counteract their biases. Hence, users are more likely to overcome some of their biases when there are more arguments in ArguCast and more opinions have been expressed by them, important implications for works considering argumentative models of human reasoning, e.g. Rago et al. (2023).
For future work, we plan to run a similar study to the crowd-sourced study with argumentation experts to determine whether our notion of coherence is intuitive to them. Additionally, we aim to extend our notion of coherence to provide a degree of coherence rather than a binary outcome. This would allow us to weight predictions, giving more influence to those that are more coherent, thereby enhancing the impact on the group forecast. Finally, we would like to use FORECAST2023 Gorska et al. (2024) and/or the hybrid-forecasting competition Benjamin et al. (2023) dataset to populate an argumentation-based forecasting system (e.g. ArguCast) and study whether the notion of coherence improves forecasting accuracy, as was the case for the notion of coherencein Irwin et al. (2022).
@@ Acknowledgments
This research was partially funded by the ERC under the EU’s Horizon 2020 research and innovation programme (grant agreement no. 101020934, ADIX), by J.P. Morgan and by the Royal Academy of Engineering, UK, under the Research Chairs and Senior Research Fellowships scheme. Any views or opinions expressed herein are solely those of the authors.
@@ Appendix A Desirable Properties of Forecaster QBAFs
This property ensures that if a forecaster has not voted on any argument then the strength of the forecasting argument $f$ should be 0.50.50.5, which is the base score of the forecasting argument. So, based on Definition 2, if the forecaster has not expressed their opinion then their prediction should be close to $\xi_{2}(f)$. This will force forecasters to vote (i.e., show their reasoning) if they want to make a prediction on $f$ further away from $\xi_{2}(f)$.
@@
We need to show that in forecaster QBAFs, if $\mathcal{V}(a)=?$ $\forall a\in\mathcal{X}$ then $\forall f\in\mathcal{F}$ $\sigma(f)=0.5$ holds. It can be seen, by Definition 1, that if $a\not\in\mathcal{F}$ and $\mathcal{V}_{u}(a)=?$ then $\tau(a)=0$, and if $a\in\mathcal{F}$ then $\tau(a)=0.5$. Since the forecaster QBAF is a tree, we first consider some $a$ which is a leaf, i.e., $\nexists b\in\mathcal{X}$ such that $(b,a)\in\mathcal{A}\cup\mathcal{S}$, by the definition of DF-QuAD, we know that $\sigma(a)=\tau(a)$. Thus, for any $f\in\mathcal{F}$ which is a leaf, the property holds. If $\exists f\in\mathcal{F}$ which are not leaves, we calculate the strengths of the arguments which are only attacked or supported by leaves, and as shown by the fact that DF-QuAD satisfies Amgoud and Ben-Naim [2018] Principle 6 (neutrality), as well as the fact that, by Definition 1, (×$(\mathcal{F}\times\mathcal{X})\cap(\mathcal{A}\cup\mathcal{S})=\emptyset$, we know that for these arguments also, $\sigma(a)=\tau(a)$ holds by the definition of DF-QuAD. We then continue to propagate up the tree in the same way, until $\forall f\in\mathcal{F}$, it is shown that $\sigma(f)=0.5$, and thus the property holds.
@@
If the forecaster agrees with only the attackers (or supporters), the final strength $\sigma(f)$ should be less (or more) than 0.50.50.5. This differs from the notion in Gorur et al. [2023], where when there were many (more than 50) supporters (or attackers) which the forecaster had no opinion and the only argument with which the forecaster agrees is an attacker (or supporter), then the strength would be exactly $\sigma(f)=0.5$, which is the base score. This would identify forecasters who exhibit bias, as their prediction would deviate from their opinions.

### v2

@@ 6.2.2 ArguCast with ArgLLMs
In order to populate ArguCast, we used LLMs to get predictions and ArgLLMs, leveraging LLMs, to construct QBAFs (we used the same prompts and hyperparameters as in Freedman et al. (2025)): we prompted the LLMs to generate supporting and attacking arguments; for each generated argument, we prompted the LLMs again to obtain an uncertainty score (acting as the base score). We investigated two variations of ArgLLMs: generating arguments at breadth $1,1$ (only generating a supporter and an attacker for each forecasting argument), and breadth $n,k$ (generating $n$ supporting and $k$ attacking arguments for each forecasting argument). We set the base score for each forecasting argument to 0.5. Figure 3 shows an example (at breadth $1,1$).
@@
Figure 3: An example of ArgLLM (with Mistral-7B-Instruct-v0.3 as the underpinning LLM) applied to the forecasting argument $f$ at the top, giving a QBAF with a supporting and an attacking argument, with base scores 0.6 and 0.4 respectively, leading to a strength $\sigma(f)\!=\!0.6$ with DF-QuAD (given $\tau(f)\!=\!0.5$). The resulting forecaster QBAF leads to an incoherent forecast, as the 40% prediction is lower than $\xi_{2}(f)\!=\!0.5$, but $\sigma(f)\!>\!\xi_{1}(f)\!=\!0.5$.
@@ 6.3 Results
Table 2 reports the forecasting accuracy for each LLM and the forecasting accuracy in the set of coherent predictions for both breadth $1,1$ and breadth $n,k$. The results indicate that applying coherence generally leads to improvements or, at least, preserves forecasting accuracy across all models. Note that applying coherence significantly reduces the total number of predictions, at breadth $n,k$ generally retaining more samples compared to breadth $n,k$. Most notably, Llama 3 retains only 44% of samples at breadth $1,1$, while it preserves 71% at breadth $n,k$. The most samples retained were with the Mistral model at breadth $n,k$, preserving 78% of the samples.
@@
Table 2: Accuracy comparison between raw LLMs and ArgLLMs with breadth $1,1$ and breadth $n,k$. For raw LLMs, we show accuracy as (correct/total). For ArgLLMs we show the percentages of coherent and correct predictions within those of coherent predictions (coherent & correct/total coherent). The bold percentage shows the best variant for the given model. The total columns show the number of samples retained for each model.
The key finding is that accuracy consistently improves when coherence is applied, particularly at breadth $1,1$. GPT-4o shows the most dramatic improvement, with accuracy increasing from 69% to 80% at breadth $1,1$. Llama 3 and Mixtral also achieve their best performance when coherence is applied at breadth $1,1$, reaching 82% and 78% accuracy, respectively. Interestingly, Mistral is the only model that achieved a higher accuracy at breadth $n,k$ compared to breadth $1,1$, which could suggest that already better performing models may require more arguments to be generated.
These results support a positive answer to RQ2. Notably, the improvements are most consistent at breadth $1,1$. Although small, the consistency of these improvements is valuable, as in the case of RQ1.
@@ 7 Crowd-sourced Experiments
Figure 4: Example of a debate shown to users. This includes a forecasting argument (in cyan), a supporter thereof (in gray with a green solid arrow pointing towards the forecasting argument), an attacker of the forecasting argument (also in gray but with a red dotted arrow pointing towards the forecasting argument), and the votes of a fictitious forecaster $u$ (where blue thumbs up represents $V(u,a)=+$ and pink thumbs down represents $V(u,a)=$-).
@@ 7.1 Research Questions
We anticipated, for RQ3, the null hypothesis (stating that users’ reasoning and our notion of coherence disagree in the same manner) not to be rejected. However, for RQ4, we expected the null hypothesis to be rejected (i.e., that RQ4 is answered affirmatively).
@@
$|\mathcal{F}|=1\land|\mathcal{D}|=2\land|\mathcal{A}|=1\land|\mathcal{S}|=1\land(a,f)\in\mathcal{A}\rightarrow$
$\mathcal{V}(u,a)=+\land(s,f)\in\mathcal{S}\rightarrow\mathcal{V}(u,s)=-.$
@@
$(\exists_{c\in\mathcal{D}}\exists_{d\in\mathcal{D}}\exists_{g\in\mathcal{F}}(c,g)\in\mathcal{A}\land(d,g)\in\mathcal{S}\land\mathcal{V}(u,c)=\mathcal{V}(u,d)).$
@@
A breadth complex variant is an ACF $\langle\mathcal{X}=\mathcal{F}\cup\mathcal{D},\mathcal{R}=\mathcal{A}\cup\mathcal{S},\mathcal{U},\mathcal{V},\mathcal{P}\rangle$, such that:
$\lvert\{a\in\mathcal{D}|\not\exists_{r\in\mathcal{R}}{(r=(a_{1},a),a_{1}\in\mathcal{D})}\}\rvert=3$.
@@
$|\{a\in\mathcal{D}|\exists_{r\in\mathcal{R}}{(r=(a,a_{1}),a_{1}\in\mathcal{X})}\}|=1$.
A depth complex variant occurs when the forecaster adds a new argument to any of the leaf (i.e. unattacked and unsupported) nodes. This argument can either support or attack one of the child arguments (the supporter or attacker of the forecasting argument).11
@@ 7.2 Experimental Design
In this experiment, we set $\xi_{1}(f)=\xi_{2}(f)=0.5$. We choose $\xi_{2}(f)=0.5$ due to the forecasting questions being fictitious and having a 50% probability of occurring. The participants were divided into two groups: 50% of them were asked to assess simpler questions first and more complex questions later (named the Simple-Complex group) while the other 50% of were asked to assess more complex questions first and simpler questions after (the Complex-Simple group). In both cases, users were randomly allocated (50%, 50%) which (of the two) forecasting questions to assess. For the four (variants of) debates the users saw, Alex’s forecast was also allocated randomly (1/3, 1/3, 1/3) amongst $50\%,=50\%,>50\%$. In Appendix E we further discuss the experimental design of this study.
@@ 7.4 Results
Out of 100 participants, only 46 passed all attention checks and were included in our study. For RQ3 we analysed 184 samples using McNemar’s test McNemar (1947) used for determining whether two classification models disagree in the same way. The null hypothesis was that the two marginal probabilities for each outcome (from our notion and users) were identical (i.e., H0: $12=76$, H1: $12\neq 76$ from the contingency Table 3). The chi-squared test (commonly used in conjunction with McNemar’s) with one degree of freedom yielded $p<.00001$. Given the result is significant at $p<0.05$, we rejected H0 in favour of H1, indicating that the users’ opinion of coherence does not align with our notion.
@@
Table 3: Contingency table for McNemar’s test. Columns/rows represent the number of debate variants the users/our notion of coherence (respectively) found coherent for the same debate variants.
@@
[…]
Table 4: Number of (not) aligned and pairs (alignment between our notion of coherence and the participants in the experiment) for the debate variations: simple, breadth complex, depth complex and vote complex.
Table 5 summarises the results for RQ4.1-RQ4.3. For each complexity type, we conducted a two-sample one sided t-test to determine if complex debate variants aligned more with our notion of coherence compared to the simpler variants. Our null hypothesis for all RQ4.1-RQ4.3 was H0: $\mu_{\text{complex}}\leq\mu_{\neg\text{complex}}$ against the alternative H1: $\mu_{\text{complex}}>\mu_{\neg\text{complex}}$. All three complexity types were statistically significant at $p<0.05$. Vote complexity yielded a p-value of 0.009, breadth complexity a p-value of 0.04, and depth complexity the most significant effect with a p-value of 0.002. Based on these results, we can reject all three null hypotheses.
@@
Table 5: Mean and standard deviation of aligned pairs (between our notion of coherence and the participants from the user experiment) for each complex and not-complex debate variations. These values are derived from Table 4. Then, it reports on the p-value derived from the two-sample one sided t-test.
Rejecting RQ3 suggests that the notion of coherence with $\xi_{1}(f)=\xi_{2}(f)=0.5$ does not accurately reflect the intuition of what users perceive as coherent. This further implies the need for a coherence constraint, as users are not naturally coherent in their judgment. Additionally, since we have rejected RQ4.1, RQ4.2, and RQ4.3 and these were refinements of RQ4, we can also reject RQ4. This means that our notion of coherence aligns more with users’ reasoning when debates are complex. This could suggest that our notion of coherence is more intuitive when more information is available to the users. Furthermore, this indicates that with complex debate variations, the users spend more cognitive effort with complex debate variations, which helps them counteract their biases. We have left the qualitative error analysis of forecaster biases to future work.
@@ 8 Conclusion
In this paper, we introduced a novel notion of argumentative coherence for judgmental forecasting, formally defined in terms of alignment between forecasters’ internal argumentative reasoning and their predictions. Then, through crowd-sourced and in-person user studies, as well as experiments with LLMs, we systematically evaluated the usefulness of the coherence property.
First, in an in-person user study with ArguCast, we found that filtering out incoherent forecasts improved the aggregated predictions, moving them closer to expert baselines. Second, in the experiments with LLMs, we applied coherence to LLM-generated forecasts and ArgLLM-generated argumentative reasoning and demonstrated consistent improvements in accuracy across different models. These results from the in-person and LLM experiments support the practical need for integrating argumentative coherence into judgmental forecasting systems (ArguCast, GJOpen, Metaculus). Finally, in the crowd-sourced experiments, we found that the users’ opinions may not align with (our notion of) coherence. This suggests that some form of coherence constraints are necessary. Additionally, we observe that the users’ opinions aligned more with our notion of coherence when debates were more complex. We speculate that this is due to our notion being more intuitive when more information is available. The results also indicated that as users spend more cognitive effort with complex debates (which implies that they reason more), it helps them counteract their biases. This may have implications for works considering argumentative models of reasoning, e.g. Rago et al. (2023).
For future work, we plan to run a similar study to the crowd-sourced study with argumentation experts to determine whether our notion of coherence is intuitive to them. Additionally, we aim to extend our notion of coherence to provide a degree of coherence rather than a binary outcome. This would allow us to weight predictions, giving more influence to those that are more coherent, thereby enhancing the impact on the group forecast. We also plan to explore relationships between our approach and probabilistic argumentation Toni et al. (2023). Finally, we would like to use FORECAST2023 Gorska et al. (2024) and/or the hybrid-forecasting competition Benjamin et al. (2023) dataset to populate an argumentation-based forecasting system (e.g. ArguCast) and study whether the notion of coherence improves forecasting accuracy, as was the case for the notion of coherence in Irwin et al. (2022).
@@ Acknowledgments
This research was partially funded by the ERC under the EU’s Horizon 2020 research and innovation programme (grant agreement no. 101020934, ADIX), by J.P. Morgan and by the Royal Academy of Engineering, UK, under the Research Chairs and Senior Research Fellowships scheme (grant agreement no. RCSRF2021\11\45).
Supplementary Material
@@ Appendix A Desirable Properties of Forecaster QBAFs
This property ensures that if a forecaster has not voted on any argument then the strength of the forecasting argument $f$ should be $0.5$, which is the base score of the forecasting argument. So, based on Definition 2, if the forecaster has not expressed their opinion then their prediction should be close to $\xi_{2}(f)$. This will force forecasters to vote (i.e., show their reasoning) if they want to make a prediction on $f$ further away from $\xi_{2}(f)$.
@@
We need to show that in forecaster QBAFs, if $\mathcal{V}(a)=?$ $\forall a\in\mathcal{X}$ then $\forall f\in\mathcal{F}$ $\sigma(f)=0.5$ holds. It can be seen, by Definition 1, that if $a\not\in\mathcal{F}$ and $\mathcal{V}_{u}(a)=?$ then $\tau(a)=0$, and if $a\in\mathcal{F}$ then $\tau(a)=0.5$. Since the forecaster QBAF is a tree, we first consider some $a$ which is a leaf, i.e., $\nexists b\in\mathcal{X}$ such that $(b,a)\in\mathcal{A}\cup\mathcal{S}$, by the definition of DF-QuAD, we know that $\sigma(a)=\tau(a)$. Thus, for any $f\in\mathcal{F}$ which is a leaf, the property holds. If $\exists f\in\mathcal{F}$ which are not leaves, we calculate the strengths of the arguments which are only attacked or supported by leaves, and as shown by the fact that DF-QuAD satisfies Amgoud and Ben-Naim [2018] Principle 6 (neutrality), as well as the fact that, by Definition 1, $(\mathcal{F}\times\mathcal{X})\cap(\mathcal{A}\cup\mathcal{S})=\emptyset$, we know that for these arguments also, $\sigma(a)=\tau(a)$ holds by the definition of DF-QuAD. We then continue to propagate up the tree in the same way, until $\forall f\in\mathcal{F}$, it is shown that $\sigma(f)=0.5$, and thus the property holds.
@@
If the forecaster agrees with only the attackers (or supporters), the final strength $\sigma(f)$ should be less (or more) than $0.5$. This differs from the notion in Gorur et al. [2023], where when there were many (more than 50) supporters (or attackers) which the forecaster had no opinion and the only argument with which the forecaster agrees is an attacker (or supporter), then the strength would be exactly $\sigma(f)=0.5$, which is the base score. This would identify forecasters who exhibit bias, as their prediction would deviate from their opinions.